OpenStudy (anonymous):

Find the square root of 2+5i so that the real part of your answer is positive. The square root is ? If I square it i get 4+25i^2 4-25 -21 Which would be -21+0i Right? But how do i make the real part positive and what is the square root?

OpenStudy (anonymous):

I guess u assume a root of a+ib, square it, then compare real and imaginary parts.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

If i do that, I get 2 = a^2-b^2 and 5 = 2ab to solve.

OpenStudy (anonymous):

I guess u could also use DeMoivre, if u know that way...

OpenStudy (anonymous):

um... not to make myself look ignorant but I dont know what you mean where did (2 = a^2-b^2 and 5 = 2ab) come from and what is and a and b?

OpenStudy (anonymous):

I am assuming there is a square root of the form a + ib so if u square that u get a^2 + 2abi -b^2 which put equal to 2 + 5i so a^2-b^2 = 2 and 2ab = 5 equating real and imaginary parts.